OpenStudy (anonymous):

Can someone Help? For a Function F(X) = B^X the Domain is: All Reals All Reals Where X > 0 All Reals where y > 0 (0, *Infinty*)

OpenStudy (allank):

I'm thinking All Reals.

OpenStudy (allank):

Let's say the function is f(x) = 2^x Try x=-1, f(x) = 2^(-1) = 1/2 That knocks out choice B and D.

OpenStudy (allank):

Try setting y=0. 0=2^x log0=xlog2 x=log0/log2 log0 is undefined, thus this knocks out C because there is no value of x that can yield y that is equal to, or less than 0.

OpenStudy (anonymous):

Dude Thanks So much! i have one more question to ask but i'll make a new one so you can get another medal :P

OpenStudy (allank):

You're welcome =)